Circular No. 7019

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 R. Remillard and E. Morgan, Massachusetts Institute of Technology;

and J. McClintock and G. Sobczak, Harvard-Smithsonian Center for Astrophysics, write: “The x-ray nova and blackhole candidate XTE J1550–564 has brightened dramatically since its discovery at 0.07 Crab (2–12 keV) on Sept. 7 (IAUC 7008). Observations with the RXTE All Sky Monitor show a steady rise to 1.7 Crab on Sept. 15. Thereafter, there is increased variability, with the intensity reaching 3.2 Crab on Sept. 18.7 UT. On Sept. 19 and 20, a large flare peaked at 6.8 Crab, and the intensity fell back to the range of 2.7–3.6 Crab on Sept. 20 and 21. PCA observations on Sept. 22.15 indicate a flux of 2.0 Crab at 2–30 keV. The long-term trend shows a steeper x-ray spectrum with increased brightness, while a local hardening of the spectrum was observed during the Sept. 19–20 flare. This x-ray nova is the brightest yet observed by RXTE, and observations at other wavelengths are strongly encouraged.”

J. D. Hawley, S. M. Petrinec, and D. L. Chenette, Lockheed Martin Missiles and Space ATC Space Physics, communicate: “XTE J1550–564 (IAUC 7008) has been observed by the PIXIE instrument onboard the NASA/POLAR spacecraft since it first became visible on Sept. 7. The first PIXIE observation with adequate counting statistics was at Sept. 7.7 UT. The object is visible for ∼ 5.5 hr every orbit (18-hr period, apogee over north pole at 9 R_{⊕}). Although the primary mission of PIXIE is to study the bremsstrahlung x-rays of the earth’s auroral regions, we also have the opportunity to observe celestial x-ray sources that appear off the limb of the earth. Averaged x-ray flux measurements (E = 2–12-keV, 64 logarithmically spaced energy channels) for each apogee pass show a nearlylinear slope of +0.8 photon s^{−1} cm^{−2} day^{−1} for Sept. 7–16. On Sept. 16, the x-ray flux dipped but then rose rapidly (flux of 5.8 photons s^{−1} cm^{−2} on Sept. 16.0, versus 24 photons s^{−1} cm^{−2} on Sept. 19.5), with the flux peaking on Sept. 20. PIXIE also measures Cir X-1, and on Sept. 12 the intensity of XTE J1550–564 surpassed the typical level of Cir X-1. Preliminary analysis shows similar spectral shapes for Cir X-1 and XTE J1550–564. Spectra along with more detailed information and plots can be found on

the PIXIE home page, http://pixie.spasci.com.”

SUPERNOVA 1998eb IN NGC 1961 S. Pesci, Milan, Italy, reports that a visual search of NGC 1961 with a

0.5-m telescope on Aug. 26.9 UT showed no new star (limiting mag 15.8).

1998 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green